# FocalCodec: Low-Bitrate Speech Coding via Focal Modulation Networks

### v1

@@ D.1 FocalCodec
The compressor processes 1024-dimensional WavLM features and forwards them through 3 focal downscaling blocks with hidden dimensions of 1024, 512, and 256, respectively. Each block has two focal levels, a window size of 7, a focal factor of 2, and a layer scale initialization of 0.0001. A final projection maps the 256-dimensional hidden states to latent representations of dimension 13, which are then quantized with a binary spherical codebook of 213 $=$ 8192 codes. The decompressor mirrors the compressor, replacing focal downscaling blocks with focal upscaling blocks to reconstruct the 1024-dimensional continuous representations from the quantized latent codes. We use a weight of 1.0 for the reconstruction loss and a weight of 0.1 for the entropy loss. We train on LibriTTS (Zen et al., 2019) (585 hours from 2456 speakers) using full utterances rather than fixed-length chunks, which differs from related work. This approach allows us to fully exploit the unlimited receptive field of focal modulation. This is in line with our vision that the encoder should be as powerful as possible to extract high-quality representations, while the decoder can be lightweight and use limited context windows. For this stage, we use the AdamW (Loshchilov & Hutter, 2019) optimizer with an initial learning rate of 0.0005, $\beta_{1}$ of 0.8, $\beta_{2}$ of 0.99, and weight decay of 0.01. The learning rate is reduced by a factor of 0.9 if validation loss does not improve within a margin of 0.0025. Gradients are clipped to a maximum L2 norm of 5. Training stops when validation loss does not decrease for several consecutive epochs.
@@
The decoder processes 1024-dimensional WavLM features and forwards them through 8 ConvNeXt blocks with a hidden dimension of 512, a feed-forward dimension of 1536, a kernel size of 7, and padding of 3. For the STFT, we set the FFT size to 1024 samples and the hop length to 320. The feature matching loss is calculated using 80-dimensional log-Mel spectrograms with the same STFT configuration. The discriminator adopts the convolutional architecture introduced in (Kong et al., 2020). We train on LibriTTS using audio chunks of 7040 samples with a batch size of 16. Due to resource constraints, our training is limited to the train-clean-100 split. We found this amount of data sufficient to obtain high-quality reconstructions. We use the AdamW optimizer with an initial learning rate of 0.0002, $\beta_{1}$ of 0.8, $\beta_{2}$ of 0.99, and weight decay of 0.01. The learning rate follows an exponential decay schedule with a factor of 0.999. Training continues until perceived audio quality stops improving, which occurs around 3M steps.
@@
For the smaller variant of FocalCodec used in the ablation studies, we employ the same setup with the following modifications: the hidden sizes in the three focal downscaling blocks are reduced from 1024, 512, 256 to 512, 256, 128; the codebook size is decreased to 1024; the model is trained on LibriSpeech train-clean-100 using a batch size of 4.
D.2 Downstream Tasks
@@
The model architecture is a 2-layer BiLSTM with 512-dimensional hidden states. A CTC (Graves et al., 2006) head is stacked on top and trained to predict either characters or BPE units. Experiments use characters and BPE vocabularies of sizes 250, 500, and 1000, with the best result reported. Note that for Mimi and FocalCodec@12.5, training on characters is infeasible due to the low token rate (12.5 Hz), which results in hidden sequences shorter than the target, making them incompatible with CTC loss. For all models except Mimi, performance improves monotonically with increasing BPE sizes up to 1000, while Mimi achieves the best results with BPE-500. If the codec employs multiple codebooks, we compute a weighted sum of the embeddings from each codebook, with the weights learned during training, as done in (Chen et al., 2022). The embedding layer is initialized using the discrete embeddings from the codec quantizer.
@@
The model architecture is a Conformer (Gulati et al., 2020) encoder with 6 layers, 4 attention heads, a model dimension of 512, and a feed-forward layer dimension of 2048. Codecs with multiple codebooks use a weighted sum of embeddings for the input, with independent linear heads for each codebook in the output. The embedding layer is initialized using the discrete embeddings from the codec quantizer. Training is performed using cross-entropy loss between predicted and target tokens.
@@
The SS setup closely mirrors that of SE. The only difference is that training is performed using cross-entropy loss with permutation invariant training (Kolbæk et al., 2017), and the number of output heads is doubled to account for predicting two sources in parallel.
@@
The model architecture is an autoregressive Llama 3 (Grattafiori et al., 2024) decoder with 12 layers, 4 attention heads, 1 key-value head, a model dimension of 512, a feed-forward layer dimension of 2048, and a base RoPE frequency of 10,000. To provide speaker information, we extract speaker embeddings from the target utterance using WavLM-base (Chen et al., 2022), fine-tuned for speaker verification. The pooled speaker embedding is prepended to the text embeddings to condition the model on speaker identity. The embedding layer is initialized using the discrete embeddings from the codec quantizer. Training is performed with next-token prediction, where the input sequence consists of pooled speaker embedding, text embeddings, and speech token embeddings. The cross-entropy loss is computed only on speech tokens, while the text and speaker embeddings are excluded from loss computation. For inference, we use top-$p$ sampling with $p=\text{0.9}$ and temperature of 1.0.
@@
For all tasks, we use AdamW (Loshchilov & Hutter, 2019) optimizer with a batch size of 16, an initial learning rate of 0.0001, $\beta_{1}$ = 0.8, $\beta_{2}$ = 0.99, weight decay of 0.01, and dropout of 0.1. The learning rate is reduced by a factor of 0.9 if validation loss does not improve within a margin of 0.0025. Gradients are clipped to a maximum L2 norm of 0.01. Training stops if validation loss does not decrease for several consecutive epochs.
Appendix E Implementation and Hardware
Software for the experimental evaluation was implemented in Python using the SpeechBrain (Ravanelli et al., 2021, 2024) toolkit. Each model is trained on a single GPU, with the choice between V100 GPUs (16 or 32 GB) and A100 GPUs (40 GB), depending on cluster resource availability.
Appendix F Additional Results
F.1 Subjective Evaluation
We conduct a subjective test with 21 participants who rate a total of 10 reconstructions from LibriSpeech test-clean. Following prior work (Défossez et al., 2023; Zhang et al., 2024; Liu et al., 2024; Parker et al., 2024), we employ the MUSHRA (Schoeffler et al., 2018) format without hidden anchor. Listeners compare multiple versions of an example at once, including a labeled reference and a hidden reference. They are asked the following question: “Please evaluate the quality proximity between an audio sample and its reference. Please listen carefully to the reference audio and then rate the quality of each test audio clip compared to the reference. Use the scale where 0 indicates no resemblance to the reference, and 100 means perfectly the same as the reference.” Participants were recruited online by sharing a link to the test across various public channels. To keep the subjective test short, we selected a subset of baselines based on their overall performance in objective metrics [^6] All the baselines are presented on the project page to foster reproducibility and fair comparison.. As showcased in Figure 2, FocalCodec can achieve extremely low bitrates without compromising performance significantly. We observe that FocalCodec is Pareto-optimal until 0.50 kbps. Its performance is superior to most baselines and remains comparable to BigCodec, Stable Codec and WavTokenizer. Finally, we emphasize that FocalCodec strikes the optimal trade-off between MOS and downstream task performance, balancing perceived audio quality and objective metrics.
Figure 2: Subjective evaluation from 21 participants averaged over 10 samples. Left. Trade-off between mean opinion score and bitrate. The green dashed line highlights the reference score. Right. Distribution of mean opinion score. The red lines highlight the median. FocalCodec@50 median is marginally lower than BigCodec, Stable Codec and WavTokenizer. However, user preference remains comparable when accounting for variability.
F.2 Comparison to TS3-Codec
TS3-Codec (Wu et al., 2024a) is a recent transformer-only architecture designed for low-bitrate streaming speech coding. Despite its lower bitrate and streamable architecture, it remains competitive with BigCodec, the current state-of-the-art. Like FocalCodec, it utilizes a single quantizer. However, its fully transformer-based architecture prioritizes reconstruction, focusing on acoustic representations. The model was trained on Libri-Light (Kahn et al., 2020). Since the model is not publicly available, we reached out to the authors to obtain reconstructions of the LibriSpeech test-clean for comparison. Table 8 shows the results. FocalCodec@50 surpasses TS3-Codec across all evaluated metrics, while FocalCodec@25, despite operating at a significantly lower bitrate, still achieves superior performance in terms of UTMOS and dWER. These findings further highlight the effectiveness of the proposed models.
Table 8: Clean speech resynthesis on LibriSpeech test-clean.
@@ F.2 Comparison to TS3-Codec
F.3 Chunk-Wise Streaming Inference
@@
Although our codec is non-causal, it can be made streamable via chunked inference. This involves splitting the input signal into fixed-size chunks with a certain amount of overlap to reduce boundary artifacts. To assess the streamability of our codec, we use a left context of 48,000 samples (3 seconds) and an overlap of 250 samples while varying the chunk size. The reconstructed chunks are stitched together using the overlap–add method with linear fade-in/fade-out. As shown in Table 9, FocalCodec@50 maintains acceptable performance when streamed with a chunk size of 8000 (500 milliseconds). In contrast, for FocalCodec@25 and FocalCodec@12.5, performance drops significantly. This is because, due to the 2x and 4x downsampling along the time dimension, they require a much larger look-ahead.
Table 9: Offline vs chunk-wise streaming inference.
Codec Chunk Size UTMOS $\uparrow$ dWER $\downarrow$ Sim $\uparrow$
LibriSpeech test-clean
FocalCodec@50 Inf 4.05 2.18 97.4
FocalCodec@25 Inf 4.14 3.30 96.3
FocalCodec@12.5 Inf 4.22 7.94 93.9
FocalCodec@50 2000 (125 ms) 2.17 6.06 95.9
FocalCodec@50 4000 (250 ms) 2.71 4.62 96.6
FocalCodec@50 8000 (500 ms) 3.16 4.55 96.9
FocalCodec@25 8000 (500 ms) 2.95 12.17 95.6
FocalCodec@12.5 8000 (500 ms) 2.84 47.43 91.8
F.4 Speech Separation with Continuous Input
Although our codec outperforms the baselines in speech separation (see Section 4.5), its performance remains insufficient for practical use. We hypothesize that this limitation stems not from the lack of expressive power of the continuous representations – since WavLM is pretrained on simulated noisy and overlapping speech, its lower layers retain the ability to represent mixtures (Chen et al., 2022) – but rather from the fact that our compressor, quantizer, decompressor, and decoder are not explicitly trained on mixtures. As a result, the quantized representations struggle to disentangle speakers. To investigate this further, we replaced the discrete input representations with continuous representations from the encoder. We focused on FocalCodec@50 and WavLM-KM6, the top two models in terms of intelligibility of reconstructed speech, both of which share the same encoder (WavLM-large layer-6). As shown in Table 10, using continuous representations as input instead of discrete ones significantly improves performance, with FocalCodec@50 outperforming WavLM-KM6 in both dWER and Sim. This suggests that WavLM-large layer-6 retains substantial information relevant to speech separation. Training the quantizer with mixtures could further improve the quality of the learned tokens for this task.
Table 10: Discrete vs continuous input features for speech separation.
Codec Input Features DNSMOS $\uparrow$ dWER $\downarrow$ Sim $\uparrow$
Libri2Mix
Reference 3.77 0.00 100.0
WavLM6-KM Discrete 3.49 76.91 85.0
WavLM6-KM Continuous 3.68 23.09 89.4
FocalCodec@50 Discrete 3.71 73.87 89.0
FocalCodec@50 Continuous 3.76 17.35 93.8
F.5 Mel-Spectrogram Analysis
@@
Figure 3 shows examples of reconstructed Mel-spectrograms from LibriSpeech (left) and Libri1Mix (right), using the 3 top-performing codecs. The reconstructed speech from LibriSpeech is almost indistinguishable from the ground truth. For Libri1Mix, the first row shows audio contaminated with noise, while the second row shows the original clean audio. It can be observed that BigCodec, a purely acoustic codec trained for reconstruction, attempts to reconstruct the noise, resulting in poor intelligibility. In contrast, Stable Codec and FocalCodec, which have semantically meaningful representations, are able to perform basic denoising. Notably, FocalCodec assigns more energy to the frequency bands corresponding to speech, even more than in the original clean audio, leading to improved intelligibility. On the other hand, Stable Codec, while providing good denoising, introduces some artifacts and static noise in the lower part of the spectrogram, which degrades intelligibility.
@@
Figure 3: Reconstructed Mel-spectrograms from LibriSpeech (left) and Libri1Mix (right).

### v2

@@ E.1 FocalCodec
The compressor processes 1024-dimensional WavLM features and forwards them through 3 focal downscaling blocks with hidden dimensions of 1024, 512, and 256, respectively. Each block has two focal levels, a window size of 7, a focal factor of 2, and a layer scale initialization of 0.0001. A final projection maps the 256-dimensional hidden states to latent representations of dimension 13, which are then quantized with a binary spherical codebook of 213 = 8192 codes. The decompressor mirrors the compressor, replacing focal downscaling blocks with focal upscaling blocks to reconstruct the 1024-dimensional continuous representations from the quantized latent codes. We use a weight of 1.0 for the reconstruction loss and a weight of 0.1 for the entropy loss. We train on LibriTTS [84] (585 hours from 2456 speakers) using full utterances rather than fixed-length chunks, which differs from related work. This approach allows us to fully exploit the unlimited receptive field of focal modulation. This is in line with our vision that the encoder should be as powerful as possible to extract high-quality representations, while the decoder can be lightweight and use limited context windows. For this stage, we use the AdamW [40] optimizer with an initial learning rate of 0.0005, $\beta_{1}$ of 0.8, $\beta_{2}$ of 0.99, and weight decay of 0.01. The learning rate is reduced by a factor of 0.9 if validation loss does not improve within a margin of 0.0025. Gradients are clipped to a maximum L2 norm of 5. Training stops when validation loss does not decrease for several consecutive epochs.
@@
The decoder processes 1024-dimensional WavLM features and forwards them through 8 ConvNeXt blocks with a hidden dimension of 512, a feed-forward dimension of 1536, a kernel size of 7, and padding of 3. For the STFT, we set the FFT size to 1024 samples and the hop length to 320. The feature matching loss is calculated using 80-dimensional log-Mel spectrograms with the same STFT configuration. The discriminator adopts the convolutional architecture introduced in [32]. We train on LibriTTS using audio chunks of 7040 samples with a batch size of 16. Due to resource constraints, our training is limited to the train-clean-100 split. We found this amount of data sufficient to obtain high-quality reconstructions. We use the AdamW optimizer with an initial learning rate of 0.0002, $\beta_{1}$ of 0.8, $\beta_{2}$ of 0.99, and weight decay of 0.01. The learning rate follows an exponential decay schedule with a factor of 0.999. Training continues until perceived audio quality stops improving, which occurs around 3M steps.
@@
For the smaller variant of FocalCodec used in the ablation studies, we employ the same setup with the following modifications: the hidden sizes in the three focal downscaling blocks are reduced from 1024, 512, 256 to 512, 256, 128; the codebook size is decreased to 1024; we use HiFi-GAN-V1 [32] decoder instead of Vocos [62]; the model is trained on LibriSpeech train-clean-100 using a batch size of 4.
E.2 Downstream Tasks
@@
The model architecture is a 2-layer BiLSTM with 512-dimensional hidden states. A CTC [18] head is stacked on top and trained to predict either characters or BPE units. Experiments use characters and BPE vocabularies of sizes 250, 500, and 1000, with the best result reported. Note that for Mimi and FocalCodec@12.5, training on characters is infeasible due to the low token rate (12.5 Hz), which results in hidden sequences shorter than the target, making them incompatible with CTC loss. For all models except Mimi, performance improves monotonically with increasing BPE sizes up to 1000, while Mimi achieves the best results with BPE-500. If the codec employs multiple codebooks, we compute a weighted sum of the embeddings from each codebook, with the weights learned during training, as done in [9]. The embedding layer is initialized using the discrete embeddings from the codec quantizer.
@@
The model architecture is a Conformer [19] encoder with 6 layers, 4 attention heads, a model dimension of 512, and a feed-forward layer dimension of 2048. Codecs with multiple codebooks use a weighted sum of embeddings for the input, with independent linear heads for each codebook in the output. The embedding layer is initialized using the discrete embeddings from the codec quantizer. Training is performed using cross-entropy loss between predicted and target tokens.
@@
The SS setup closely mirrors that of SE. The only difference is that training is performed using cross-entropy loss with permutation invariant training [31], and the number of output heads is doubled to account for predicting two sources in parallel.
@@
The model architecture is an autoregressive Llama 3 [17] decoder with 12 layers, 4 attention heads, 1 key-value head, a model dimension of 512, a feed-forward layer dimension of 2048, and a base RoPE frequency of 10,000. To provide speaker information, we extract speaker embeddings from the target utterance using WavLM-base [9], fine-tuned for speaker verification. The pooled speaker embedding is prepended to the text embeddings to condition the model on speaker identity. The embedding layer is initialized using the discrete embeddings from the codec quantizer. Training is performed with next-token prediction, where the input sequence consists of pooled speaker embedding, text embeddings, and speech token embeddings. The cross-entropy loss is computed only on speech tokens, while the text and speaker embeddings are excluded from loss computation. For inference, we use top-$p$ sampling with $p$ = 0.9 and temperature of 1.0. Following the experimental protocol of [65], we generated 5 samples per utterance and selected the one with the lowest WER relative to the input text, using Whisper-small [53] to obtain the transcription.
@@
For all tasks, we use AdamW [40] optimizer with a batch size of 16, an initial learning rate of 0.0001, $\beta_{1}$ = 0.8, $\beta_{2}$ = 0.99, weight decay of 0.01, and dropout of 0.1. The learning rate is reduced by a factor of 0.9 if validation loss does not improve within a margin of 0.0025. Gradients are clipped to a maximum L2 norm of 0.01. Training stops if validation loss does not decrease for several consecutive epochs.
Appendix F Implementation and Hardware
Software for the experimental evaluation was implemented in Python using the SpeechBrain [55, 54] toolkit. Each model is trained on a single GPU, with the choice between V100 GPUs (16 or 32 GB) and A100 GPUs (40 GB), depending on cluster resource availability.
Appendix G Additional Results
G.1 Comparison to TS3-Codec
TS3-Codec [74] is a recent transformer-only architecture designed for low-bitrate streaming speech coding. Despite its lower bitrate and streamable architecture, it remains competitive with BigCodec, the current state-of-the-art. Like FocalCodec, it utilizes a single quantizer. However, its fully transformer-based architecture prioritizes reconstruction, focusing on acoustic representations. The model was trained on Libri-Light [29]. Since the model is not publicly available, we reached out to the authors to obtain reconstructions of the LibriSpeech test-clean for comparison. Table˜7 shows the results. FocalCodec@50 surpasses TS3-Codec across all evaluated metrics, while FocalCodec@25, despite operating at a significantly lower bitrate, still achieves superior performance in terms of UTMOS and dWER. These findings further highlight the effectiveness of the proposed models.
Table 7: Clean speech resynthesis on LibriSpeech test-clean [47].
@@ G.1 Comparison to TS3-Codec
G.2 Subjective Evaluation
We conduct a subjective test with 40 participants who rate a total of 10 reconstructions from LibriSpeech test-clean. Following prior work [14, 86, 37, 48], we employ the MUSHRA [60] format without hidden anchor. Listeners compare multiple versions of an example at once, including a labeled reference and a hidden reference. They are asked the following question: “Please evaluate the quality proximity between an audio sample and its reference. Please listen carefully to the reference audio and then rate the quality of each test audio clip compared to the reference. Use the scale where 0 indicates no resemblance to the reference, and 100 means perfectly the same as the reference." Participants were recruited online by sharing a link to the test across various public channels. To keep the subjective test short, we selected a subset of baselines based on their overall performance in objective metrics. To ensure that participants spent sufficient time on each listening task, we filtered out submissions where less than 60 seconds were spent on any of the 10 reconstructions. Out of 40 total submissions, this resulted in 33 valid entries. As showcased in Figure˜2, FocalCodec achieves extremely low bitrates while maintaining strong performance. In particular, FocalCodec@50 outperforms most baselines and remains comparable to BigCodec and Stable Codec.
Figure 2: Subjective evaluation from 33 participants averaged over 10 samples. Left. Trade-off between mean opinion score and bitrate. The green dashed line highlights the reference score. FocalCodec achieves extremely low bitrates while maintaining strong performance. Right. Distribution of mean opinion score. The red lines highlight the mean. FocalCodec@50 outperforms most baselines and remains comparable to BigCodec and Stable Codec.
G.3 Streaming Inference
@@
Although our codec is non-causal, it can be made streamable via chunked inference. This involves splitting the input signal into fixed-size chunks with a certain amount of overlap to reduce boundary artifacts. To assess the streamability of our codec, we use a chunk size of 500 milliseconds with a $\sim$4% overlap and a left context of 3 seconds. The reconstructed chunks are stitched together using the overlap-add method with linear fade-in/fade-out. As shown in Table˜8 (upper section), despite the gap with the full-context model, FocalCodec@50 maintains acceptable performance at 500 milliseconds latency. However, this is still too high for strict real-time use. To enable real-time streaming, we replace all non-causal components with their causal counterparts, and we distill the non-causal WavLM features into the causal model using a feature-matching loss during training. Additionally, we train the decoder to reconstruct 24 kHz waveforms. As shown in Table˜8 (lower section), with these adjustments, together with increased model capacity and additional training data, we can achieve competitive performance at 80 milliseconds latency while maintaining a real-time factor suitable for deployment on consumer-grade GPUs.
Table 8: Streaming clean speech resynthesis on LibriSpeech test-clean [47].
Codec Bitrate (kbps) Sample Rate (kHz) Token Rate (Hz) Codebooks Causal Latency (ms) UTMOS $\uparrow$ dWER $\downarrow$ Sim $\uparrow$
FocalCodec @50 0.65 16 50.0 1 $\times$ 8192 Inf 4.05 2.18 97.4
FocalCodec @50 0.65 16 50.0 1 $\times$ 8192 500 3.16 4.55 96.9
EnCodec 1.50 24 75.0 2 $\times$ 1024 20 1.58 8.08 93.8
Mimi 0.69 24 12.5 5 $\times$ 2048 80 3.29 5.73 96.0
FocalCodec @50 0.60 24 50.0 1 $\times$ 4096 80 3.87 4.38 96.3
G.4 Mel-Spectrogram Analysis
@@
Figure˜3 shows examples of reconstructed Mel-spectrograms from LibriSpeech [47] (left) and Libri1Mix [12] (right), using the 3 top-performing codecs. The reconstructed speech from LibriSpeech is almost indistinguishable from the ground truth. For Libri1Mix, the first row shows audio contaminated with noise, while the second row shows the original clean audio. It can be observed that BigCodec, a purely acoustic codec trained for reconstruction, attempts to reconstruct the noise, resulting in poor intelligibility. In contrast, Stable Codec and FocalCodec, which have semantically meaningful representations, are able to perform basic denoising. Notably, FocalCodec assigns more energy to the frequency bands corresponding to speech, even more than in the original clean audio, leading to improved intelligibility. On the other hand, Stable Codec, while providing good denoising, introduces some artifacts and static noise in the lower part of the spectrogram, which degrades intelligibility.
@@
Figure 3: Reconstructed Mel-spectrograms from LibriSpeech [47] (left) and Libri1Mix [12] (right).